OpenStudy (anonymous):

open parentheses 5 plus square root of 3 close parentheses open parentheses 5 minus square root of 3 close parentheses

OpenStudy (anonymous):

Please explain the steps too please ?

OpenStudy (anonymous):

\[(5+\sqrt{3})(5-\sqrt{3})\]

OpenStudy (anonymous):

difference of squares \[a^2-b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

or you could FOIL it and still get the same thing

OpenStudy (anonymous):

kinda confused

OpenStudy (anonymous):

we fill in a^2 - b^2 with what ?